# Homogeneous potentials, Lagrange’s identity and Poisson geometry

### v3

@@ 3 Second family of potentials obtained
instead of the Lagrange condition (1.2). Let us compare a particular solution of Euler’s equation (1.3),
@@
where $f$ is an arbitrary smooth function of projective coordinates, with a similar particular solution of equation (3.1), which has the form
@@
If the function $U(q)$ satisfies the equation (3.1) bivector
$\bar{P}^{\prime}=$ $\left(\frac{a_{j}p_{i}}{m_{i}}-\frac{a_{i}p_{j}}{m_{j}}\right)\frac{\partial}{\partial q_{i}}\wedge\frac{\partial}{\partial q_{j}}+\left(\frac{k}{2m_{i}}p_{i}p_{j}+a_{i}\frac{\partial U}{\partial q_{j}}\right)\frac{\partial}{\partial q_{i}}\wedge\frac{\partial}{\partial p_{j}}$
$+$ $\frac{k}{2}\left(p_{i}\frac{\partial U}{\partial q_{j}}-p_{j}\frac{\partial U}{\partial q_{i}}\right)\frac{\partial}{\partial p_{i}}\wedge\frac{\partial}{\partial p_{j}}\,$, (3.3)
is a Poisson bivector invariant with respect to the phase flow of the vector field $X$
$\mathcal{L}_{X}\bar{P}^{\prime}=0\,,\qquad X=\sum_{i=1}^{n}\left(\frac{p_{i}}{m_{i}}\frac{\partial}{\partial q_{i}}-\frac{\partial U}{\partial q_{i}}\dfrac{\partial}{\partial p_{i}}\right)\,.$
Proof: The Lie derivative of the tensor field $\bar{P}^{\prime}$ (3.3) along the vector field $X$ is zero
$\mathcal{L}_{X}\bar{P}^{\prime}=\frac{p_{i}}{m_{i}}\frac{\partial\bar{Z}}{\partial q_{j}}\frac{\partial}{\partial q_{i}}\wedge\frac{\partial}{\partial p_{j}}-\left(\frac{\partial\bar{Z}}{\partial q_{i}}\frac{\partial U}{\partial q_{j}}-\frac{\partial\bar{Z}}{\partial q_{j}}\frac{\partial U}{\partial q_{i}}\right)\frac{\partial}{\partial p_{i}}\wedge\frac{\partial}{\partial p_{j}}=0\,$,
since $\bar{Z}=0$ (3.1). Because of this, the tensor field $\bar{P}^{\prime}$ is a tensor invariant of $X$.
The verification of the Jacobi identity (1.10) is reduced to checking the Jacobi conditions for the Poisson brackets between variables $x=(q,p)$, which in this case take the form
$\{q_{i},q_{j}\}^{\prime}=$ $\left(\frac{a_{j}p_{i}}{m_{i}}-\frac{a_{i}p_{j}}{m_{j}}\right)\,,\qquad\{q_{i},p_{j}\}^{\prime}=\frac{k}{2m_{i}s}p_{i}p_{j}+a_{i}\frac{\partial V}{\partial q_{j}}$
$\{p_{i},p_{j}\}^{\prime}=$ $\frac{k}{2}\left(p_{i}\frac{\partial V}{\partial q_{j}}-p_{j}\frac{\partial V}{\partial q_{i}}\right)\,.$
For example, let’s write out one of the conditions
$\{\{q_{i},q_{j}\},p_{k}\}+\{\{q_{j},p_{k}\},q_{i}\}+\{\{p_{k},q_{i}\},q_{j}\}=-\left(\frac{a_{j}p_{i}}{m_{i}}-\frac{a_{i}p_{j}}{m_{j}}\right)\dfrac{\partial}{\partial q_{k}}\bar{Z}=0\,$,
where $i,j,k=1,\ldots,n$, which is true since $\bar{Z}=0$ (3.1). The Jacobi conditions for the Poisson brackets between the remaining coordinates $x=(q,p)$ are similarly verified.
[…]
$\bar{P}^{\prime}dH=kHX\,$,
but, as above, in the space of invariant bivectors, we can find a second non-degenerate Poisson bivector in addition to $P$ (1.6).
@@
If $k\neq 0$, then the tensor field
$\tilde{P}=-\frac{k}{2}HP+\hat{P}^{\prime}\,,\qquad\mbox{rank}\tilde{P}=n\,$, (3.4)
is a non-degenerate invariant Poisson bivector, and the tensor field
$\tilde{\omega}=\tilde{P}^{-1}$
is a symplectic form invariant with respect to the phase flow generated by the Hamiltonian vector field $X$.
The proof consists of a direct verification of non-degeneracy, invariance, and Jacobi identities
$\mathcal{L}_{X}\tilde{P}=0\,,\qquad[\![\tilde{P},\tilde{P}]\!]=0\qquad\mbox{and}\qquad\mathcal{L}_{X}\tilde{\omega}=0\,,\qquad d\tilde{\omega}=0$
which is based on the fulfillment of the condition $\bar{Z}=0$ (3.1).
According to the Lie-Darboux theorem, there exists a canonical transformation of variables $(q~,p~)\phi$$\to(\tilde{q},\tilde{p})$ such that
$\tilde{P}=\sum_{i=1}^{n}\frac{\partial}{\partial\tilde{q}_{i}}\wedge\frac{\partial}{\partial\tilde{p}_{i}}\qquad\mbox{and}\qquad\tilde{\omega}=\sum_{i=1}^{n}d\tilde{p}_{i}\wedge d\tilde{q_{i}}\,$.
@@
The Poisson bivector $\tilde{P}$ satisfies the relation
$\tilde{P}dH=\frac{k}{2}\,H\,PdH$
which includes two invariant non-degenerate Poisson tensors and can be considered as a tensor analogue of the identity (3.2) for inhomogeneous potentials.
@@ 4 Conclusion
In this paper we prove an existence of additional tensor invariants for non-integrable Hamiltonian systems satisfying the Lagrange identity on cotangent bundle $T^{*}\mathbb{R}^{n}$. We also show similar invariant Poisson bivectors for the systems on the Lie-Poisson algebras $so^{*}(4)$ and $e^{*}(3)$.

### v4

@@ 3 Nonhomogeneous potentials of a special kind
The well-known particular solution of Euler’s equation (1.4) has the following form
@@
where $f$ is an arbitrary smooth function of projective coordinates. Similar particular solution of equation (3.1), has the form
@@
For the vector field
$\bar{Y}=\sum_{i=1}^{n}a_{i}\frac{\partial}{\partial q_{i}}+\frac{k}{2}\sum_{i=1}^{n}p_{i}\frac{\partial}{\partial p_{i}}\,$, (3.3)
our Hamiltonian $H=T+U$ is the Darboux invariant at $\bar{Z}=0$ since
$\bar{Y}(T+U)=k(T+U)+Pd\bar{Z}\,.$
The corresponding vector field
$\bar{X}=PdH=\left(\sum_{i=1}^{n}\frac{p_{i}}{m_{i}}\frac{\partial}{\partial q_{i}}\right)-\left(\sum_{j=1}^{n}\frac{\partial U}{\partial q_{j}}\dfrac{\partial}{\partial p_{j}}\right)$
is also a Darboux invariant for the vector field $\bar{Y}$ (3.3)
$\bar{Y},\bar{X}]=\left(\frac{k}{2}-1\right)\bar{X}\qquad\mbox{at}\qquad\bar{Z}=0\,.$
From this, the following statement immediately follows.
@@
Second-rank bivector
$\bar{P}^{\prime}=\bar{X}\wedge\bar{Y}$
is the invariant Poisson bivector for the phase flow of the vector field $\bar{X}$ if the function $U(q)$ satisfies equation (3.1), that is, if
$\bar{Y}(H)=kH\qquad\mbox{at}\qquad H\neq 0\,.$
The proof consists of substituting the Lie bracket between the vector fields $\bar{X}$ and $\bar{Y}$ into the equations
$\mathcal{L}_{\bar{X}}\,\bar{P}^{\prime}=[\bar{X},\bar{X}]\wedge\bar{Y}+\bar{X}\wedge[\bar{X},\bar{Y}]=0\,.$
and
$\bar{P}^{\prime},\bar{P}^{\prime}]\!]=2\bar{X}\wedge[\bar{Y},\bar{X}]\wedge\bar{Y}=0\,.$
In terms of the variables $q$ and $p$, this Poisson bivector takes the following form
$\bar{P}^{\prime}=$ $\sum_{1\leq i<j\leq n}\left(\frac{a_{j}p_{i}}{m_{i}}-\frac{a_{i}p_{j}}{m_{j}}\right)\frac{\partial}{\partial q_{i}}\wedge\frac{\partial}{\partial q_{j}}+\sum_{i=1}^{n}\sum_{j=1}^{n}\left(\frac{k}{2m_{i}}p_{i}p_{j}+a_{i}\frac{\partial U}{\partial q_{j}}\right)\frac{\partial}{\partial q_{i}}\wedge\frac{\partial}{\partial p_{j}}$
$+$ $\sum_{1\leq i<j\leq n}\frac{k}{2}\left(p_{i}\frac{\partial U}{\partial q_{j}}-p_{j}\frac{\partial U}{\partial q_{i}}\right)\frac{\partial}{\partial p_{i}}\wedge\frac{\partial}{\partial p_{j}}\,.$ (3.4)
The Poisson bivector $\bar{P}^{\prime}$ (3.4) is degenerate for $n\geq 2$, since $\bar{P}^{\prime}$ has rank 2, and satisfies the relation
$P^{\prime}dH=kHX\,.$
In the general case, the Poisson bivector $\bar{P}^{\prime}$ (3.4) is not commutative with the Poisson bivector P{P} (1.7). Nevertheless, as in the case of homogeneous potentials, there exists a unique linear combination of these invariant bivectors with coefficients depending on the Hamiltonian $H$ and the parameter $k$, which is the invariant Poisson tensor.
@@
If the potential $U(q)$ satisfies the relation (3.1) for $k\neq 0$, then on the open subset defined by the inequality $H\neq 0$, the tensor field of valence $(2,0)$
$\bar{P}=-\frac{k}{2}HP+\bar{P}^{\prime}\,,\qquad\mbox{rank}\bar{P}=n\,$, (3.5)
is a non-degenerate invariant Poisson bivector, and the $(0,2)$-valence tensor field
$\bar{\omega}=\bar{P}^{-1}$
is a symplectic form invariant under the phase flow generated by the Hamiltonian vector field $\bar{X}$.
The proof consists of a direct verification.
According to the Lie-Darboux theorem, there exists a canonical transformation of variables $\phi:(q,p)\to(\bar{q},\bar{p})$ such that
$\bar{P}=\sum_{i=1}^{n}\frac{\partial}{\partial\bar{q}_{i}}\wedge\frac{\partial}{\partial\bar{p}_{i}}\qquad\mbox{and}\qquad\bar{\omega}=\sum_{i=1}^{n}d\bar{p}_{i}\wedge d\bar{q_{i}}\,.$
@@
For all three Poisson invariant vectors $P$, $\bar{P}^{\prime}$, and $\bar{P}$, the Hamiltonian fluxes generated by the Hamiltonian function $H$ with potential $U(q)$ (3.1) are proportional to
$\bar{P}^{\prime}dH=kH\bar{X}\qquad\mbox{and}\qquad\bar{P}dH=\frac{k}{2}H\bar{X}$
where $\bar{X}=PdH$, i.e., they can be reduced to each other by a time substitution.
@@ 4 Conclusion
This paper proves the existence of additional tensor invariants for non-integrable Hamiltonian systems, provided that the Hamiltonian function $H$ is a Darboux invariant with respect to the Euler type vector fields, which is equivalent to the condition that the Lagrangian identity or its analogue holds.
The existence of such tensor invariants for non-integrable systems raises a number of questions about the properties of these invariant geometric structures, about the qualitative behavior of trajectories that preserve these geometric invariants, about what distinguishes these non-integrable systems from the rest, about the use of these invariants for numerical integration, and so on. We do not yet know the answers to these questions.
The author thanks A.J. Maciejewski and an anonymous reviewer for their extensive comments on the first version of this work, which allowed for its significant improvement